### v1

@@ V-B Training on All Perturbations ablation
We report results on training PerAct with 100 demos (with batch size 16 for 300k iterations) in All Perturbations setting in Figure 7. Zero-shot evaluation of a PerAct model trained on RLBench variations data with no perturbation_factors enabled achieves task-averaged success rate of 6.4% (28.1% lower than No Perturbations task-averaged success rate). When we train the model with with All Perturbations enabled, the task-averaged success rate increases by 21.1% only. However, the model should be able to perform the same tasks under any environmental setting for being practically deployable. This result indicates that The Colosseum’s perturbation_factors not only study systematic perturbations added to the environment, but also increase the difficulty of the tasks itself, even with ground truth perturbed scenes available for training.
@@ V-C Real-world alignment analysis for The Colosseum
We first observed only a marginal difference in success rate of 6.67% during evaluation on the No Perturbations tasks between PerAct trained in simulation and that in real-world settings. This served as a crucial sanity check for tasks performance between the two models before advancing forward to evaluating them across the 12 perturbation_factors. We observed that for factors such as MO_Texture, Light_Color, Table_Color, Table_Texture, and Distractor, the discrepancies in performance between both models on each individual factor were marginal, remaining under 5%.
@@ VII Conclusion
We introduced The Colosseum, a comprehensive benchmark designed to assess the generalization capabilities of Behavior Cloning (BC) models in robotic manipulation. The Colosseum systematically perturbs the task environments of the robot’s workspace along an exhaustive list of axes — including object appearance and size, lighting, background, table-top appearance, and camera pose — both in simulation and real-world. Through empirical studies conducted with SotA BC methods on The Colosseum, we identified which perturbation factors most significantly impact model’s success rates on tasks they are trained to execute. Additionally, we demonstrated a close alignment between The Colosseum in simulated and real-world. To enhance reproducibility and facilitate future model evaluations in both simulated and real-world, we will open-source our resources along with the 3D printed assets. The Colosseum offers a platform for future research to develop and quantitatively evaluate robotic manipulation models before scaling via a unified leaderboard.

### v2

@@ V-B Training on All Perturbations ablation
We report results on training PerAct with 100 demos (with batch size 16 for 300k iterations) in All Perturbations setting in Figure 7. Zero-shot evaluation of a PerAct model trained on RLBench variations data with no perturbation_factors enabled achieves task-averaged success rate of 6.4% (28.1% lower than No Perturbations task-averaged success rate). When we train the model with with All Perturbations enabled, the task-averaged success rate increases by 21.1% only. However, the model should be able to perform the same tasks under any environmental setting for being practically deployable. This result indicates that The Colosseum’s perturbation_factors not only study systematic perturbations added to the environment, but also increase the difficulty of the tasks itself, even with ground truth perturbed scenes available for training. From this ablation with All Perturbations as proxy for an extreme case of factor compounding, which is results in 28.1% lower in success rates with respect to No Perturbations, hence suggesting that compounding perturbation factors has some degree of compounding effect toward models’ performance.
@@ V-C Real-world alignment analysis for The Colosseum
We first observed only a marginal difference in success rate of 6.67% during evaluation on the No Perturbations tasks between PerAct trained in simulation and that in real-world settings. This served as a crucial sanity check for tasks performance between the two models before advancing forward to evaluating them across the 14 perturbation_factors. We observed that for factors such as MO_Texture, Light_Color, Table_Color, Table_Texture, and Distractor, the discrepancies in performance between both models on each individual factor were marginal, remaining under 5%.
Figure 8: Example rollouts using The Colosseum perturbations in tandem with comparable real-world scenarios. We combine various The Colosseum variations to form three combinations of compounding perturbation scenarios, each paired with a curated real-world scene (workbench, dining table, study room table) featuring the same set of perturbations that are naturally present in the scene, similar to the approach used by [9]. We evaluate these combinations using PerAct model trained on No Perturbation, and aim to establish a correlation between our simplified compounding perturbation scenarios and realistic real-world scenes.
@@
V-D The Colosseum perturbations grounded in the real-world
To validate that the perturbations in The Colosseum accurately mimic naturally occurring environmental or object variations in real-world scenarios, we carried out an ablation study. This study compared three realistic scenes—workbench, dining table, and study room tabletop—with corresponding perturbation combinations derived from The Colosseum as shown in Figure 8. Utilizing a multitask instance of PerAct trained with No Perturbation. We evaluated all the pairs of scenarios for the task of slide_block_to_target. Over five trials with ten episodes each, we observed a significant correlation. Notably, the combination of [Distractor + MO_Size] resulted in an ${R}^{2}=0.75$, while the combination of [Light_Color + Table_Texture + Distractor + MO_Size] achieved an ${R}^{2}=0.83$. For further details on the results and methodology, please refer to Supplementary Section IV.E.
@@ VII Conclusion
We introduced The Colosseum, a comprehensive benchmark designed to assess the generalization capabilities of Behavior Cloning (BC) models in robotic manipulation. The Colosseum systematically perturbs the task environments of the robot’s workspace along an exhaustive list of axes — including object appearance and size, lighting, physical properties of objects, background, table-top appearance, and camera pose — both in simulation and real-world. Through empirical studies conducted with SotA BC methods on The Colosseum, we identified which perturbation factors most significantly impact model’s success rates on tasks they are trained to execute. Additionally, we demonstrated a close alignment between The Colosseum in simulated and real-world. To enhance reproducibility and facilitate future model evaluations in both simulated and real-world, we will open-source our resources along with the 3D printed assets. The Colosseum offers a platform for future research to develop and quantitatively evaluate robotic manipulation models before scaling via a unified leaderboard.
VIII Acknowledgement
Jiafei Duan is supported by the Agency for Science, Technology and Research (A*STAR) National Science Fellowship. We would also like to thank Yi Li and the members of the UW RSE Lab for their feedback on the paper.